you are planning to save for retirement over the next 15 years to do this you will invest $700 a month in a stock account and $400 a month in a bond account. The return of the stock account is expected to be 10 precent and the bond account will pay 6 precent. When you retire, you will combine your money into an account with a 8 precent return.
How much can you withdraw each month from your account assuming a 20 year withdrawal period ?
Answer #1

FV of investment in stock = annuity*[(1+r)^n -1 /r] r = 10/12 = .83% n = 15*12 =180

700*[(1.0083)^180 -1 /.0083]

FV = 700*[415.815] = 291071

FV of investment in bond market = annuity*[(1+r)^n -1 /r] r= 6/12 =.5% n = 15*12 = 180

400*[(1.005)^180 -1 /.005]

FV = 400*[290.818] = 116327.5

total fund at the time of retirement = 116327.5+291071 = 407398.5

Formula for PMT = r*PV / 1-(1+r)^-n r =8/12 = .666% n = 20*12 =240

.00666*407398.5 / 1-(1.00666)^-240

2713.274 / .7967 = 3405.616

amount to be withdrawn per month = 3405.616